# Privacy Bills of Materials: A Transparent Privacy Information Inventory for Collaborative Privacy Notice Generation in Mobile App Development

### v1

@@ Abstract
Privacy regulations mandate that developers must provide authentic and comprehensive privacy notices, e.g., privacy policies or labels, to inform users of their apps’ privacy practices. However, due to a lack of knowledge of privacy requirements, developers often struggle to create accurate privacy notices, especially for sophisticated mobile apps with complex features and in crowded development teams. To address these challenges, we introduce Privacy Bills of Materials (PriBOM), a systematic software engineering approach that leverages different development team roles to better capture and coordinate mobile app privacy information. PriBOM facilitates transparency-centric privacy documentation and specific privacy notice creation, enabling traceability and trackability of privacy practices. We present a pre-fill of PriBOM based on static analysis and privacy notice analysis techniques. We demonstrate the perceived usefulness of PriBOM through a human evaluation with 150 diverse participants. Our findings suggest that PriBOM could serve as a significant solution for providing privacy support in DevOps for mobile apps.
1 Introduction
@@
Due to functional, analytical, and advertising needs, mobile application developers are increasingly expanding their collection and use of users’ personal information and other privacy-related data. Many privacy regulations, such as the General Data Protection Regulation (GDPR) [6], the California Consumer Privacy Act (CCPA) [4], and the Australian Privacy Principles (APP) [1], require developers to provide authentic and understandable privacy notices to inform users of the app’s privacy practices. Privacy policy is the most prevailing format of privacy notices to mobile application users [21, 30, 36, 31, 65, 37, 44, 64]. In pursuit of higher readability and conciseness, mainstream application stores in the market have also recently required app developers to remind users of app’s potential privacy data practices in the form of privacy nutrition labels (a.k.a. privacy label) [41, 42, 43], e.g. the Data Safety Sections (DSS) in Google Play and the Apple Privacy Labels (APL) in Apple’s App Store.
@@
However, numerous studies [24, 25, 52, 61] have shown that existing app privacy notices are often problematic, as they fail to authentically align with the actual data collection and usage behaviors from apps. While non-comprehensive and inaccurate privacy notices could harm users’ trustworthiness and violate privacy regulations, software developers face various challenges when providing authoritative privacy notices. Crafting a good privacy notice is complex, requiring not only legal knowledge but also a fundamental understanding of the app’s various functions and features. Developers commonly lack training or knowledge in privacy and legal fields [52, 49, 50], and they often hold a passive or even negative attitude towards privacy factors during development [52, 50, 48]. Additionally, even though legal teams are mainly responsible for the privacy notice documentation, they are naturally not acquainted with mobile app technical details. Such numerous challenges have promoted the development of assistance tools in generating privacy notices.
@@
There are three types of mainstream tools to help the development team generate privacy notices: Online Automated Privacy Policy Generators (APPGs) [63, 7, 3, 17, 13, 14], Code-based Privacy Policy Generators (CPPGs) [86, 85, 90] and the recently emerged Code-based IDE Plugins (CIDEPs) [49, 51, 53]. Although these tools are useful and widely adopted, most are not applicable to sophisticated mobile apps with complex features and in crowded development teams. A small change in a basic function could require significant effort to accurately reflect the privacy practices in privacy notices. Concurrently, thousands of such modifications will occur in the software development and maintenance process, making it impossible to manually track all those privacy practice changes, ultimately leading to problematic privacy notices. These issues are further magnified in the development under multi-role collaboration. The legal team often feel like “lone wolves” carrying the company’s privacy program alone [9], given that other roles are minimally involved in managing privacy. Thus, existing generators are far from enough, a systematic and collaborative software engineering solution involving various roles is pressingly needed.
@@
Focusing on facilitating collaborative privacy notice generation in mobile app development, PriBOM is a table-like privacy information inventory indexed by UI widget, documenting information regarding 1) UI Identifier, 2) Codebase and Permission, 3) Third-Party Library and 4) Privacy Notice Disclosure. Table 2 presents the format of PriBOM. The UI widget serves as both visual elements and key components in functionality and data handling, therefore it is the pivot to synergistically connect different roles of developers on privacy-related communication. Additionally, we present a pre-fill of PriBOM based on cutting-edge static analysis and privacy notice analysis techniques, demonstrating its practicability. Furthermore, we conducted a usability evaluation of PriBOM through a survey of 150 participants. The statements about design intuitiveness, content comprehension, and information relevance of PriBOM receive positive feedback, underlined by the 85.3%, 72%, and 78.76% agreement, respectively. We observed differences in the perspectives held by different roles, with experience within the same role further shaping these viewpoints. Additionally, we summarized the most frequently mentioned aspects of PriBOM as identified by non-technical roles, with the efficiency benefit widely recognized for its ability to streamline workflows. Lastly, we discussed potential enhancements and ongoing refinements to ensure the practical adaptation of PriBOM and its alignment with real-world demands across various roles.
@@
  • We conduct a human evaluation to comprehensively assess the usefulness of PriBOM.
Figure 1: Use cases of PriBOM. (1) A privacy information inventory indexed by UI widgets, providing transparent privacy documentation. (2) A privacy communication platform between different roles in the development team. (3) A systematic solution for collaborative privacy notice generation.
2 Privacy Notice Generation for Mobile Apps
2.1 Status Quo
Although privacy policies have become the primary privacy notice approach for mobile applications [21, 30, 36, 31, 65, 37, 44], their presentation and readability have always been criticized [57, 28]. To improve the usability, Kelly et al. [41, 42, 43] introduced the privacy nutrition labels, or privacy labels, designed to facilitate consumers’ understanding of how their information is collected and utilized in a concise and structured manner. Privacy labels have been widely adopted by practitioners and have become a trend for conveying apps’ privacy practices to end users. As required by privacy regulations (e.g., GDPR), providing accurate privacy notices is equally important as providing them in an inviting way. Developers are responsible for creating accurate privacy notices [53], though various challenges and concerns have been discovered and raised [52, 24, 25], even for big companies such as Google [^1] In re Facebook, Inc. Internet Tracking Litigation, 956 F. 3d 589 - Court of Appeals, 9th Circuit, 2020 and Facebook [^2] In re Google Assistant Privacy Litigation, 457 F. Supp. 3d 797 - Dist. Court, ND California. Failing to do so may cause serious legal consequences. To respond to those challenges, prior efforts have been made to assist developers create privacy notices. Based on their inherent nature, these tools can be categorized into various groups, including Online Automated Privacy Policy Generators (APPGs) [63, 7, 3, 17, 13, 14], Code-based Privacy Policy Generators (CPPGs) [86, 85, 90] and Code-based IDE Plugins (CIDEPs) [49, 51, 53]. We introduce the evolution process of these tools and summarize their key features below.
@@ 2.1 Status Quo
(c) Matcha, a CIDEP proposed by Li et al. [53]
Figure 2: (a) and (b) are the interfaces of [3], one of the most popular APPGs, according to Pan et al. [63], on collecting Personally Identifiable Information (PII) and Third-party Libraries (TPL) usages. (c) is a screenshot of Matcha [53], one of the Code-based IDE Plugins (CIDEP), from its JetBrains plugin page [11].
Most APPGs are questionnaire-based [63] online tools that depend on developer-provided information to generate privacy policies, as shown in Figure 2(a) and Figure 2(b). Although APPGs can directly generate privacy policies, their qualities can be largely affected by developers’ unartistic design flaws and inaccurate completion. As the privacy policy generation process by APPGs is completely disconnected from the original software development in DevOps, it is inevitable that developers cannot appropriately maneuver these tools.
CPPGs have been proposed to further make the privacy notice generation stage “shift left” in DevOps, i.e., in the earlier stages of the software development cycle. Tools like AutoPPG [85, 86] and PrivacyFlash [90] analyze privacy-related features contained in the source code of Android and iOS applications and generate privacy descriptions or notices based on code features. Additionally, Si et al. [71] proposed Repo2Label, a framework to automatically generate privacy labels based for generative AI-based applications based on code repositories. However, CPPGs face significant challenges, such as inherent complexity and low explainability, which hinder their adoption and effectiveness. Also, CPPGs often fail to ensure compliance with high-level privacy regulations, particularly concerning non-functional requirements [63].
Unlike generating complete privacy notices post-development, CIDEPs are integrated into the integrated development environment (IDE) to provide code privacy annotations for developers during the development process. Figure 2(c) shows an example [53] of such CIDEPs. These tools, such as Coconut [49], Honeysuckle [51] and Matcha [53], help developers add privacy annotations to provide information for privacy notice creation, thereby reducing common misunderstandings among developers and easing the creation of privacy notices.
Table 1: Summary of developer’s privacy challenges and comparison of assistant tools along those challenges. "APPG" stands for Questionnaire-based Online Automated Privacy Policy Generators, "CPPG" stands for Code-based Privacy Policy Generators, "CIDEP" stands for Code-based IDE Plugins. ●: addressed. ◐: partially addressed. ○: not addressed.
@@
Misunderstanding of privacy terms [45, 52, 49, 50, 25, 73]
Lack of knowledge in privacy and legal field [45, 52, 49, 50, 25, 48, 77, 39]
Update and iteration of privacy rules [45, 50]
@@
Opacity of third-party libraries and resources [45, 52, 49, 50, 25, 70]
Complicated privacy notice creation process [45, 52]
Lack of awareness of privacy-preserving alternative implementations [49, 50, 25]
Limited tool support in understanding data practices [49, 50, 48, 74, 73]
@@
Not well-maintained privacy documentations [49]
Negative and demotivated attitude towards privacy [52, 49, 50, 25, 48, 70, 77, 39, 74]
Lack of team, organization and platform support [52, 49, 48, 70, 77, 73]
@@
2.2 Formative Study
Developers often encounter privacy challenges as they attempt to build a thorough understanding of the privacy practices in the application and convert them into privacy notices [53]. We first summarize the previous studies to analyze developers’ challenges relevant to privacy aspects when they are developing applications or creating privacy notices.
@@ 2.2 Formative Study
[Challenge-1] Privacy Knowledge Absence. Such challenges pertain to the understanding of privacy within the development community. Developers often lack privacy knowledge [45, 52, 49, 50, 25, 48, 77, 39] due to the missing of formal privacy training, e.g. interpretation of privacy-related terminology, and may not be aware of alternative approaches that offer better privacy preservation [49, 50, 25]. Misunderstanding Privacy terms [45, 52, 49, 50, 25, 73] can lead to discrepancies between intended and implemented privacy practices, potentially causing privacy issues.
@@
[Challenge-2] Limited Technical Knowledge. This kind of challenge involves privacy issues in the technical aspect, especially related to estimating a thorough understanding of the data practices. For example, developers often integrate Third-Party Libraries (TPLs) without a full understanding of their data handling practices [45, 52, 49, 50, 25, 70], leading to a lack of transparency regarding functionality and privacy of external sources. Such opacity may cause developers to struggle to follow data minimization, which is a principle introduced by privacy regulations such as GDPR [6] and formally defined by researchers [89]. Data minimization requires developers to only collect necessary personal data in relation to specific purposes. Although existing tools [49, 50, 48, 74, 73] can provide support, such necessity checks still rely on manual checks by legal experts [89, 75].
@@
[Challenge-3] Unfriendly Organizational Environment. These challenges reflect the broader environmental and motivational factors that influence privacy. The challenge of developers’ negative and demotivated attitude towards privacy is mentioned by most studies [52, 49, 50, 25, 48, 70, 77, 39, 74] and highlight the necessity of an urgent change of the attitude. There are mainly three reasons that contribute to this phenomenon. First, as privacy is considered to be a non-functional factor, trade-offs exist between privacy and other objectives like functionalities, user experience, business goals and model performance [48]. Second, the ownership of Privacy is uncertain. Developers may regard privacy as the responsibility of other personnel and not consider it in their own workflow [48]. Last, developers may find them in a workplace culture that may not prioritize privacy [74]. The absence of a supportive infrastructure for privacy within organizations can demotivate developers from pursuing stringent privacy standards. According to  [77], less than a quarter of developers have access to security experts. A disconnect between individual developers and decision-makers in the organization may also lead to negativeness towards privacy [48].

### v2

# Privacy Bills of Materials (PriBOM): A Transparent Privacy Information Inventory for Collaborative Privacy Notice Generation in Mobile App Development

@@ Abstract.
Privacy regulations mandate that developers must provide authentic and comprehensive privacy notices, e.g., privacy policies or labels, to inform users of their apps’ privacy practices. However, due to a lack of knowledge of privacy requirements, developers often struggle to create accurate privacy notices, especially for sophisticated mobile apps with complex features and in crowded development teams. To address these challenges, we introduce PriBOM (Privacy Bills of Materials), a systematic software engineering approach that leverages different development team roles to better capture and coordinate mobile app privacy information. PriBOM facilitates transparency-centric privacy documentation and specific privacy notice creation, enabling traceability and trackability of privacy practices. We present a pre-fill of PriBOM based on static analysis and privacy notice analysis techniques. We explore the perceived usefulness of PriBOM through a human evaluation with 150 diverse participants. The role of PriBOM in enhancing privacy-related communication is well received with 83.33% agreement, suggesting that PriBOM could serve as a significant solution for providing privacy support in DevOps for mobile apps.
Transparency, Usable Privacy, Mobile Applications, Privacy Policy, Privacy Paradox
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
@@
Due to functional, analytical, and advertising needs, mobile application developers are increasingly expanding their collection and use of users’ personal information and other privacy-related data. Many privacy regulations, such as the General Data Protection Regulation (GDPR) (GDP, 2016), the California Consumer Privacy Act (CCPA) (CCP, 2018), and the Australian Privacy Principles (APP) (APP, 2014), require developers to provide authentic and understandable privacy notices to inform users of the app’s privacy practices. Privacy policy is the most prevailing format of privacy notices to mobile application users (Adams, 2020; Bui et al., 2023; Flavián and Guinalíu, 2006; Caramujo and Da Silva, 2015; Perez et al., 2018; Harkous et al., 2018; Kemp, 2020). In pursuit of higher readability and conciseness, mainstream application stores in the market have also recently required app developers to remind users of app’s potential privacy data practices in the form of privacy nutrition labels (a.k.a. privacy label) (Kelley et al., 2009, 2010, 2013), e.g. the Data Safety Sections (DSS) in Google Play and the Apple Privacy Labels (APL) in Apple’s App Store.
@@
However, numerous studies (Balebako and Cranor, 2014; Balebako et al., 2014; Li et al., 2022; Pan et al., 2023) have shown that existing app privacy notices are often problematic, as they fail to authentically align with the actual data practices of apps. While non-comprehensive and inaccurate privacy notices could harm users’ trustworthiness and violate privacy regulations, software developers face various challenges when providing authoritative privacy notices. Crafting a good privacy notice is complex, requiring not only legal knowledge but also a fundamental understanding of the app’s various functions and features. Developers commonly lack training or knowledge in privacy and legal fields (Li et al., 2022, 2018, 2021a), and often hold a passive or even negative attitude towards privacy factors during development (Li et al., 2022, 2021a; Lee et al., 2024). Additionally, even though legal teams are mainly responsible for the privacy notice documentation, they are naturally not acquainted with mobile app technical details. Such numerous challenges have promoted the development of assistance tools in generating privacy notices.
@@
There are three types of mainstream tools to help the development team generate privacy notices: Online Automated Privacy Policy Generators (APPGs) (Pan et al., 2024b; Iub, 2022; App, 2024; Ter, 2024; Pri, 2024b, a), Code-based Privacy Policy Generators (CPPGs) (Yu et al., 2018, 2015; Zimmeck et al., 2021) and the recently emerged Code-based IDE Plugins (CIDEPs) (Li et al., 2018, 2021b, 2024). Although these tools are useful and widely adopted, most are not applicable to sophisticated mobile apps with complex features and in crowded development teams. A small change in a basic function could require significant effort to accurately reflect the privacy practices in privacy notices. Concurrently, thousands of such modifications will occur in the software development and maintenance process, making it impossible to manually track all those privacy practice changes, ultimately leading to problematic privacy notices. These issues are further magnified in the development under multi-role collaboration. The legal team often feel like “lone wolves” carrying the company’s privacy program alone (Lin, 2024a), given that other roles are minimally involved in managing privacy. Thus, existing generators are far from enough, a systematic and collaborative software engineering solution involving various roles is pressingly needed.
@@
Focusing on facilitating collaborative privacy notice generation in mobile app development, PriBOM is a table-like privacy information inventory indexed by UI widget, documenting information regarding 1) UI Identifier, 2) Codebase and Permission, 3) Third-Party Library and 4) Privacy Notice Disclosure. Table 2 presents the format of PriBOM. The UI widget serves as both visual elements and key components in functionality and data handling, therefore it is the pivot to synergistically connect different roles of developers on privacy-related communication. Additionally, we present a pre-fill of PriBOM based on cutting-edge static analysis and privacy notice analysis techniques, demonstrating its practicability. Furthermore, we conducted a usability evaluation of PriBOM through a survey of 150 participants. By using a survey, we also aim to prompt discussions and opinions around potential adaption to specific applications beyond only validating its usefulness. The statements about design intuitiveness, content comprehension, and information relevance of PriBOM receive positive feedback, underlined by the 85.3%, 72%, and 78.76% agreement, respectively. We observed differences in the perspectives held by different roles, with experience within the same role further shaping these viewpoints. Non-technical roles frequently highlighted PriBOM’s efficiency in streamlining workflows. Lastly, we discussed potential enhancements and ongoing refinements to ensure its practical adaptation and its alignment with real-world demands across roles. The implementation and survey questionnaire are available at: https://anonymous.4open.science/r/PriBOM-CE36.
@@
  • We conduct a human evaluation to comprehsnsively assess the usefulness of PriBOM.
Figure 1. Use cases of PriBOM. (1) A privacy information inventory indexed by UI widgets, providing transparent privacy documentation. (2) A privacy communication platform between different roles in the development team. (3) A systematic solution for collaborative privacy notice generation.
2. Privacy Notice Generation for Mobile Apps
2.1. Status Quo
Although privacy policies have become the primary privacy notice approach for mobile applications (Adams, 2020; Bui et al., 2023; Flavián and Guinalíu, 2006; Caramujo and Da Silva, 2015; Perez et al., 2018; Harkous et al., 2018; Kemp, 2020), their presentation and readability have always been criticized (McDonald and Cranor, 2008; Blakkarly and Graham, 2024). To improve the usability, Kelly et al. (Kelley et al., 2009, 2010, 2013) introduced the privacy nutrition labels, or privacy labels, designed to facilitate consumers’ understanding of how their information is collected and utilized in a concise and structured manner. Privacy labels have been widely adopted by practitioners and have become a trend for conveying apps’ privacy practices to end users. As required by privacy regulations (e.g., GDPR), providing accurate privacy notices is equally important as providing them in an inviting way. Developers are responsible for creating accurate privacy notices (Li et al., 2024), though various challenges and concerns have been discovered and raised (Li et al., 2022; Balebako and Cranor, 2014; Balebako et al., 2014), even for big companies such as Google [^1] In re Facebook, Inc. Internet Tracking Litigation, 956 F. 3d 589 - Court of Appeals, 9th Circuit, 2020 and Facebook [^2] In re Google Assistant Privacy Litigation, 457 F. Supp. 3d 797 - Dist. Court, ND California. Failing to do so may cause serious legal consequences. To respond to those challenges, prior efforts have been made to assist developers create privacy notices. Based on their inherent nature, these tools can be categorized into various groups, including Online Automated Privacy Policy Generators (APPGs) (Pan et al., 2024b; Iub, 2022; App, 2024; Ter, 2024; Pri, 2024b, a), Code-based Privacy Policy Generators (CPPGs) (Yu et al., 2018, 2015; Zimmeck et al., 2021) and Code-based IDE Plugins (CIDEPs) (Li et al., 2018, 2021b, 2024). We introduce the evolution process of these tools and summarize their key features below.
@@ 2.1. Status Quo
(c) Matcha, a CIDEP proposed by Li et al. (Li et al., 2024)
Figure 2. (a) and (b) are the interfaces of (App, 2024), one of the most popular APPGs, according to Pan et al. (Pan et al., 2024b), on collecting Personally Identifiable Information (PII) and Third-party Libraries (TPL) usages. (c) is a screenshot of Matcha (Li et al., 2024), one of the Code-based IDE Plugins (CIDEP), from its JetBrains plugin page (Mat, 2024).
Most APPGs are questionnaire-based (Pan et al., 2024b) online tools that depend on developer-provided information to generate privacy policies, as shown in Figure 2(a) and Figure 2(b). Although APPGs can directly generate privacy policies, their qualities can be largely affected by developers’ unartistic design flaws and inaccurate completion. As the privacy policy generation process by APPGs is completely disconnected from the original software development in DevOps, it is inevitable that developers cannot appropriately maneuver these tools.
CPPGs have been proposed to further make the privacy notice generation stage “shift left” in DevOps, i.e., in the earlier stages of the software development cycle. Tools like AutoPPG (Yu et al., 2015, 2018) and PrivacyFlash (Zimmeck et al., 2021) analyze privacy-related features contained in the source code of Android and iOS applications and generate privacy descriptions or notices based on code features. However, CPPGs face significant challenges, such as inherent complexity and low explainability, which hinder their adoption and effectiveness. Also, CPPGs often fail to ensure compliance with high-level privacy regulations, particularly concerning non-functional requirements (Pan et al., 2024b).
Unlike generating complete privacy notices post-development, CIDEPs are integrated into the integrated development environment (IDE) to provide code privacy annotations for developers during the development process. Figure 2(c) shows an example (Li et al., 2024) of such CIDEPs. These tools, such as Coconut (Li et al., 2018), Honeysuckle (Li et al., 2021b) and Matcha (Li et al., 2024), help developers add privacy annotations to provide information for privacy notice creation, thereby reducing common misunderstandings among developers and easing the creation of privacy notices.
Table 1. Summary of developer’s privacy challenges and comparison of assistant tools along those challenges. ”APPG” stands for Questionnaire-based Online Automated Privacy Policy Generators, ”CPPG” stands for Code-based Privacy Policy Generators, ”CIDEP” stands for Code-based IDE Plugins. ●: addressed. ◐: partially addressed. ○: not addressed.
@@
Misunderstanding of privacy terms (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Tahaei et al., 2020)
Lack of knowledge in privacy and legal field (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Lee et al., 2024; Weir et al., 2020; Keküllüoğlu and Acar, 2023)
Update and iteration of privacy rules (Khandelwal et al., 2023; Li et al., 2021a)
@@
Opacity of third-party libraries and resources (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Seymour et al., 2023)
Complicated privacy notice creation process (Khandelwal et al., 2023; Li et al., 2022)
Lack of awareness of privacy-preserving alternative implementations (Li et al., 2018, 2021a; Balebako et al., 2014)
Limited tool support in understanding data practices (Li et al., 2018, 2021a; Lee et al., 2024; Tahaei et al., 2021, 2020)
@@
Not well-maintained privacy documentations (Li et al., 2018)
Negative and demotivated attitude towards privacy (Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Lee et al., 2024; Seymour et al., 2023; Weir et al., 2020; Keküllüoğlu and Acar, 2023; Tahaei et al., 2021)
Lack of team, organization and platform support (Li et al., 2022, 2018; Lee et al., 2024; Seymour et al., 2023; Weir et al., 2020; Tahaei et al., 2020)
@@
2.2. Formative Study
Developers often encounter privacy challenges as they attempt to build a thorough understanding of the privacy practices in the application and convert them into privacy notices (Li et al., 2024). We first summarize the previous studies to analyze developers’ challenges relevant to privacy aspects when they are developing applications or creating privacy notices.
@@ 2.2. Formative Study
[Challenge-1] Privacy Knowledge Absence. Such challenges pertain to the understanding of privacy within the development community. Developers often lack privacy knowledge (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Lee et al., 2024; Weir et al., 2020; Keküllüoğlu and Acar, 2023) due to the missing of formal privacy training, e.g. interpretation of privacy-related terminology, and may not be aware of alternative approaches that offer better privacy preservation (Li et al., 2018, 2021a; Balebako et al., 2014). Misunderstanding Privacy terms (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Tahaei et al., 2020) can lead to discrepancies between intended and implemented privacy practices, potentially causing privacy issues.
@@
[Challenge-2] Limited Technical Knowledge. This kind of challenge involves privacy issues in the technical aspect, especially related to estimating a thorough understanding of the data practices. For example, developers often integrate Third-Party Libraries (TPLs) without a full understanding of their data handling practices (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Seymour et al., 2023), leading to a lack of transparency regarding functionality and privacy of external sources. Such opacity may cause developers to struggle to follow data minimization, which is a principle introduced by privacy regulations such as GDPR (GDP, 2016) and formally defined by researchers (Zhou et al., 2023). Data minimization requires developers to only collect necessary personal data in relation to specific purposes. Although existing tools (Li et al., 2018, 2021a; Lee et al., 2024; Tahaei et al., 2021, 2020) can provide support, such necessity checks still rely on manual checks by legal experts (Zhou et al., 2023; Wang et al., 2022).
@@
[Challenge-3] Unfriendly Organizational Environment. These challenges reflect the broader environmental and motivational factors that influence privacy. The challenge of developers’ negative and demotivated attitude towards privacy is mentioned by most studies (Li et al., 2022, 2018, 2021a; Balebako et al., 2014; Lee et al., 2024; Seymour et al., 2023; Weir et al., 2020; Keküllüoğlu and Acar, 2023; Tahaei et al., 2021) and highlight the necessity of an urgent change of the attitude. There are mainly three reasons that contribute to this phenomenon. First, as privacy is considered to be a non-functional factor, trade-offs exist between privacy and other objectives like functionalities, user experience, business goals and model performance (Lee et al., 2024). Second, the ownership of Privacy is uncertain. Developers may regard privacy as the responsibility of other personnel and not consider it in their own workflow (Lee et al., 2024). Last, developers may find them in a workplace culture that may not prioritize privacy (Tahaei et al., 2021). The absence of a supportive infrastructure for privacy within organizations can demotivate developers from pursuing stringent privacy standards. According to  (Weir et al., 2020), less than a quarter of developers have access to security experts. A disconnect between individual developers and decision-makers in the organization may also lead to negativeness towards privacy (Lee et al., 2024).